### v1

# The Rational Torsion Subgroup of $J_{0}(\mathfrak{p}^{r})$

Abstract
Let $\mathfrak{n}=\mathfrak{p}^{r}$ be a prime-power ideal of $\mathbb{F}_{q}[T]$ with $r\geq 2$. We study the rational torsion subgroup $\mathcal{T}(\mathfrak{p}^{r})$ of the Drinfeld modular Jacobian $J_{0}(\mathfrak{p}^{r})$. We prove that the prime-to-$q(q-1)$ part of $\mathcal{T}(\mathfrak{p}^{r})$ is equal to that of the rational cuspidal divisor class group $\mathcal{C}(\mathfrak{p}^{r})$ of the Drinfeld modular curve $X_{0}(\mathfrak{p}^{r})$. As we completely computed the structure of $\mathcal{C}(\mathfrak{p}^{r})$, it also determines the structure of the prime-to-$q(q-1)$ part of $\mathcal{T}(\mathfrak{p}^{r})$.
1 Introduction
For a positive integer $N$, let $J_{0}(N)$ be the Jacobian variety of the classical modular curve $X_{0}(N)$ and $\mathcal{T}(N):=J_{0}(N)(\mathbb{Q})_{\text{tors}}$ its rational torsion subgroup. By the Mordell-Weil theorem, $\mathcal{T}(N)$ is a finite abelian group. Let $\mathcal{C}(N)$ be the rational cuspidal divisor class group of $X_{0}(N)$ defined in [16, p. 280]. By a theorem of Manin and Drinfeld, we have
$\mathcal{C}(N)\subseteq\mathcal{T}(N).$
In the early 1970s, for any prime $p$, Ogg [10, Conj. 2] conjectured that $\mathcal{C}(p)=\mathcal{T}(p)$ and computed that $\mathcal{C}(p)$ is a cyclic group of order $\frac{p-1}{(p-1,12)}$. Later in 1977, Mazur [9, Thm. (1)] proved this conjecture by studying the Eisenstein ideal of the Hecke algebra of level $p$. There is a generalized Ogg’s conjecture:
Conjecture 1.1.
(still open) For any positive integer $N$,
$\mathcal{C}(N)=\mathcal{T}(N).$
In [16, Thm. 1.7], Yoo et al. completely determined the structure of $\mathcal{C}(N)$ for arbitrary positive $N$. Consequently, if the above conjecture holds true, then the structure of $\mathcal{T}(N)$ is determined as well. The most notable current achievement in this regard is as follows:
Theorem 1.2 (Yoo [17, Thm. 1.4], Mazur, Lorenzini, Ling, Ohta, Ren, et al.).
Let $N$ be a positive integer, and let $\ell$ be any odd prime whose square does not divide $N$. Suppose that $\ell\geq 5$, it holds that
$\mathcal{C}(N)_{\ell}=\mathcal{T}(N)_{\ell}.$
Furthermore, for $\ell=3$, the equality above remains valid under the additional condition that either $N$ is not divisible by $3$, or $N$ has a prime divisor congruent to $-1$ modulo $3$.
Remark 1.3.
In [17, (1.1)], in a particular case, when $N=p^{r}$ is a prime power with $r\geq 2$, it has been established that $\mathcal{C}(p^{r})_{\ell}=\mathcal{T}(p^{r})_{\ell}$ holds for any prime $\ell\nmid 2p$. The condition on $\ell$ here is less restrictive compared to the general case.
In the function field analogue, let $\mathbb{F}_{q}$ be a finite field of characteristic $p$ with $q$ elements. Let $A=\mathbb{F}_{q}[T]$ denote the polynomial ring in the indeterminate $T$ over $\mathbb{F}_{q}$, and $K=\mathbb{F}_{q}(T)$ represent the rational function field. Define $K_{\infty}=\mathbb{F}_{q}((\pi_{\infty}))$ as the completion of $K$ at the infinite place, and $\mathcal{O}_{\infty}=\mathbb{F}_{q}[[\pi_{\infty}]]$ as its ring of integers, where $\pi_{\infty}:=T^{-1}$. Let $|\cdot|=|\cdot|_{\infty}$ denote the normalized absolute value on $K_{\infty}$ with $|T|_{\infty}:=q$. Let $\mathbb{C}_{\infty}$ denote the completion of an algebraic closure of $K_{\infty}$. Define $\Omega=\mathbb{C}_{\infty}-K_{\infty}$ as the Drinfeld upper half plane. Let $G$ be the group scheme $\operatorname{GL}(2)$ over $\mathbb{F}_{q}$, with $Z$ denoting the group scheme of scalar matrices in $G$. Consider a nonzero ideal $\mathfrak{n}\lhd A$. The level-$\mathfrak{n}$ Hecke congruence subgroup of $G(A)$ is defined as
$\Gamma_{0}(\mathfrak{n}):=\left\{\begin{pmatrix}a&b\\ c&d\end{pmatrix}\in G(A)~{}\middle|~{}c\equiv 0~{}\text{mod}~{}\mathfrak{n}\right\}.$
Let $\Gamma_{0}(\mathfrak{n})$ act on $\Omega$ via linear fractional transformations. Drinfeld proved in [1] that the quotient $\Gamma_{0}(\mathfrak{n})\backslash\Omega$ is the space of $\mathbb{C}_{\infty}$-points of an affine curve $Y_{0}(\mathfrak{n})$ defined over $K$, representing a moduli space of rank-$2$ Drinfeld modules. The unique smooth projective curve over $K$ containing $Y_{0}(\mathfrak{n})$ as an open subvariety is denoted by $X_{0}(\mathfrak{n})$, termed as the Drinfeld modular curve of level $\mathfrak{n}$. Let $J_{0}(\mathfrak{n})$ denote the Jacobian variety of $X_{0}(\mathfrak{n})$ and $\mathcal{T}(\mathfrak{n}):=J_{0}(\mathfrak{n})(K)_{\text{tors}}$ its rational torsion subgroup. According to the Lang-Néron theorem, $\mathcal{T}(\mathfrak{n})$ is a finite abelian group. To investigate the structure of $\mathcal{T}(\mathfrak{n})$, we study the rational cuspidal divisor class group $\mathcal{C}(\mathfrak{n})$ of $X_{0}(\mathfrak{n})$, as defined in [8]. By Gekeler [6, Thm. 1.2], $\mathcal{C}(\mathfrak{n})$ is a finite group, leading to the inclusion
$\mathcal{C}(\mathfrak{n})\subseteq\mathcal{T}(\mathfrak{n}).$
The structure of $\mathcal{C}(\mathfrak{n})$ is known in some special cases. First, when $\mathfrak{n}=\mathfrak{p}\lhd A$ is a prime ideal, the structure of $\mathcal{C}(\mathfrak{p})$ was fully determined by Gekeler [5, Cor. 3.23]. Second, Papikian and Wei [14, Thm. 1.3 and Prop. 1.5] determined the prime-to-$(q-1)$ part of $\mathcal{C}(\mathfrak{n})$ for any square-free ideal $\mathfrak{n}\lhd A$. In particular, when $\mathfrak{n}=\mathfrak{p}\mathfrak{q}$ is square-free with two prime ideals $\mathfrak{p}$ and $\mathfrak{q}$ of $A$, the structure of $\mathcal{C}(\mathfrak{p}\mathfrak{q})$ was entirely determined in [12, Thm. 1.1]. Third, when $\mathfrak{n}=\mathfrak{p}^{r}\lhd A$ is a prime-power ideal with $r\geq 2$, the structure of $\mathcal{C}(\mathfrak{p}^{r})$ was completely determined in the following:
Theorem 1.4 ([8, Thm. 3.5]).
Fix a prime ideal $\mathfrak{p}\lhd A$ and $r\geq 2$. Then
$\mathcal{C}(\mathfrak{p}^{r})\cong\left(\bigoplus_{1\leq i\leq m}\frac{\mathbb{Z}}{|\mathfrak{p}|^{r-i}M(\mathfrak{p})\mathbb{Z}}\right)\oplus\left(\bigoplus_{m+1\leq i\leq r-2}\frac{\mathbb{Z}}{|\mathfrak{p}|^{i}M(\mathfrak{p})\mathbb{Z}}\right)\oplus\frac{\mathbb{Z}}{M(\mathfrak{p})\mathbb{Z}}\oplus\frac{\mathbb{Z}}{N(\mathfrak{p})\mathbb{Z}},$
where $m:=\lfloor\frac{r-1}{2}\rfloor$, $M(\mathfrak{p}):=\frac{|\mathfrak{p}|^{2}-1}{q^{2}-1}$, and
$N(\mathfrak{p}):=\begin{cases}\frac{|\mathfrak{p}|-1}{q^{2}-1},&\text{if $\deg(\mathfrak{p})$ is even.}\\ \frac{|\mathfrak{p}|-1}{q-1},&\text{otherwise.}\end{cases}$
Moreover, an explicit basis of $\mathcal{C}(\mathfrak{p}^{r})$ is provided in [8, Thm. 3.5].
Remark 1.5.
When $\mathfrak{n}\lhd A$ is square-free or $\deg(\mathfrak{n})=3$, all the cusps of $X_{0}(\mathfrak{n})$ are rational by [13, Lem. 3.1]. Thus, the cuspidal divisor groups “$\mathcal{C}(\mathfrak{n})$” of $J_{0}(\mathfrak{n})$ defined in [14, p. 522] and [13, p. 399] in the above cases are the same as the rational cuspidal divisor class groups of $X_{0}(\mathfrak{n})$ defined in [8]. However, this phenomenon does not universally hold. For instance, when $\mathfrak{n}=(T^{2}+T+1)^{2}$ and $q=2$, the cuspidal divisor group of $J_{0}(\mathfrak{n})$ is strictly larger than the rational cuspidal divisor class group of $X_{0}(\mathfrak{n})$ by [13, Ex. 8.8] and Thm. 1.4.
The following presents an analogue of generalized Ogg’s conjecture:
Conjecture 1.6.
For any non-zero ideal $\mathfrak{n}\lhd A$,
$\mathcal{C}(\mathfrak{n})=\mathcal{T}(\mathfrak{n}).$
This conjecture remains open, and the followings are some partial results:
  1. 1.
    Pál [11, Thm. 1.4] proved that $\mathcal{C}(\mathfrak{p})=\mathcal{T}(\mathfrak{p})$ for any prime ideal $\mathfrak{p}\lhd A$.
  2. 2.
    Papikian and Wei [13, Thm. 1.7] showed that $\mathcal{C}(\mathfrak{n})=\mathcal{T}(\mathfrak{n})$ for an ideal $\mathfrak{n}\lhd A$ generated by $T^{3}$ or $T^{2}(T-1)$.
  3. 3.
    Papikian and Wei [14, Thm. 1.3] demonstrated that if $\mathfrak{n}\lhd A$ is a square-free ideal, then $\mathcal{C}(\mathfrak{n})_{\ell}=\mathcal{T}(\mathfrak{n})_{\ell}$ for any prime $\ell$ not dividing $q(q-1)$.
In this paper, we prove the following:
Theorem 1.7 (Main Theorem).
Let $\mathfrak{n}=\mathfrak{p}^{r}\lhd A$ be a prime-power ideal with $r\geq 2$. For any prime $\ell\nmid q(q-1)$, we have
$\mathcal{C}(\mathfrak{p}^{r})_{\ell}=\mathcal{T}(\mathfrak{p}^{r})_{\ell}\cong\left(\frac{\mathbb{Z}_{\ell}}{M(\mathfrak{p})\mathbb{Z}_{\ell}}\right)^{r-1}\times\frac{\mathbb{Z}_{\ell}}{N(\mathfrak{p})\mathbb{Z}_{\ell}}.$
Remark 1.8.
Both the square-free case discussed in [14, Thm. 1.3] and the prime-power case in Main Theorem utilize similar approaches in their proofs. However, a notable difference arises in the choice of tools: the former employs the Atkin-Lehner involutions $W_{\mathfrak{m}}$ for studying Eisenstein ideals, while the latter relies on the Hecke operator $U_{\mathfrak{p}}$ for the same purpose.
We provide a brief overview of the proof for Main Theorem. Fix a prime-power ideal $\mathfrak{n}=\mathfrak{p}^{r}\lhd A$ with $r\geq 2$. The Hecke algebra $\mathbb{T}(\mathfrak{n})$ of level $\mathfrak{n}$ is defined as the $\mathbb{Z}$-algebra generated by all Hecke operators $T_{\mathfrak{q}}$, where $\mathfrak{q}\lhd A$ are primes, acting on the group $\mathcal{H}_{0}(\mathfrak{n},\mathbb{Z})$ of $\mathbb{Z}$-valued $\Gamma_{0}(\mathfrak{n})$-invariant cuspidal harmonic cochains on the Bruhat-Tits tree $\mathscr{T}$ of $\operatorname{PGL}(2,K_{\infty})$. For further details, please refer to [13, p. 385], Sec. 3, and Sec. 4. The Eisenstein ideal $\mathfrak{E}(\mathfrak{n})$ of $\mathbb{T}(\mathfrak{n})$ is the ideal generated by the elements
$\{T_{\mathfrak{q}}-|\mathfrak{q}|-1\mid\mathfrak{q}\text{ is prime},\mathfrak{q}\nmid\mathfrak{n}\}.$
Additionally, $\mathbb{T}(\mathfrak{n})$ naturally acts on the Jacobian $J_{0}(\mathfrak{n})$ of $X_{0}(\mathfrak{n})$; see Sec. 5 and [13, p. 386] for more information. For the prime divisor $\mathfrak{p}$ of $\mathfrak{n}$, we usually write $U_{\mathfrak{p}}:=T_{\mathfrak{p}}\in\mathbb{T}(\mathfrak{n})$. Consider the following exact sequences of groups:
[…]
where $\mathcal{C}(\mathfrak{p}^{r})[U_{\mathfrak{p}}]$ and $\mathcal{T}(\mathfrak{p}^{r})[U_{\mathfrak{p}}]$ denote the kernels of $U_{\mathfrak{p}}$. Considering the $\ell$-primary parts of the groups in the above diagram for a prime $\ell\nmid q(q-1)$, our objective is to demonstrate the following:
  1. 1.
    $U_{\mathfrak{p}}(\mathcal{C}(\mathfrak{p}^{r}))_{\ell}=U_{\mathfrak{p}}(\mathcal{T}(\mathfrak{p}^{r}))_{\ell}\cong\left(\frac{\mathbb{Z}_{\ell}}{M(\mathfrak{p})\mathbb{Z}_{\ell}}\right)^{r-2}\times\frac{\mathbb{Z}_{\ell}}{N(\mathfrak{p})\mathbb{Z}_{\ell}}.$
  2. 2.
    $\mathcal{C}(\mathfrak{p}^{r})_{\ell}[U_{\mathfrak{p}}]=\mathcal{T}(\mathfrak{p}^{r})_{\ell}[U_{\mathfrak{p}}]\cong\frac{\mathbb{Z}_{\ell}}{M(\mathfrak{p})\mathbb{Z}_{\ell}}.$
Assuming the aboves and employing the short five lemma, we obtain Main Theorem. The first statement is achieved through explicit computation of the $U_{\mathfrak{p}}$ action on the cusps of $X_{0}(\mathfrak{p}^{r})$ combined with induction on $r$. To be specific, we decompose $U_{\mathfrak{p}}$ into $U_{\mathfrak{p}}=\alpha^{\ast}\circ\beta_{\ast}$, where $\alpha^{\ast}:J_{0}(\mathfrak{n}/\mathfrak{p})\rightarrow J_{0}(\mathfrak{n})$ and $\beta_{\ast}:J_{0}(\mathfrak{n})\rightarrow J_{0}(\mathfrak{n}/\mathfrak{p})$ are defined in Sec. 5. In Lem. 5.3, we demonstrate that $\beta_{\ast}(\mathcal{C}(\mathfrak{p}^{r}))_{\ell}=\mathcal{C}(\mathfrak{p}^{r-1})_{\ell}$ for any prime $\ell\neq p$. Assume the induction hypothesis that $\mathcal{C}(\mathfrak{p}^{r-1})_{\ell}=\mathcal{T}(\mathfrak{p}^{r-1})_{\ell}$ for any prime $\ell\nmid q(q-1)$. Note that $\mathcal{C}(\mathfrak{p})=\mathcal{T}(\mathfrak{p})$ serves as the initial condition for the induction. Then for any prime $\ell\nmid q(q-1)$,
$U_{\mathfrak{p}}(\mathcal{C}(\mathfrak{p}^{r}))_{\ell}=\alpha^{\ast}(\mathcal{C}(\mathfrak{p}^{r-1}))_{\ell}=\alpha^{\ast}(\mathcal{T}(\mathfrak{p}^{r-1}))_{\ell}\supseteq U_{\mathfrak{p}}(\mathcal{T}(\mathfrak{p}^{r}))_{\ell}.$
Now, the result follows from the natural inclusion $U_{\mathfrak{p}}(\mathcal{C}(\mathfrak{p}^{r}))_{\ell}\subseteq U_{\mathfrak{p}}(\mathcal{T}(\mathfrak{p}^{r}))_{\ell}$.
The second statement is established through an Atkin-Lehner type result; see Thm. 6.5. To be precise, for a coefficient ring $R$ (see Def. 3.2), let $\mathcal{E}(\mathfrak{n},R)[U_{\mathfrak{p}}]$ represent the group of $R$-valued $\Gamma_{0}(\mathfrak{n})$-invariant harmonic cochains on $\mathscr{T}$ that are annihilated by both the actions of the Eisenstein ideal $\mathfrak{E}(\mathfrak{n})$ and the Hecke operator $U_{\mathfrak{p}}$. We construct an explicit element $E_{\mathfrak{n}}$ in $\mathcal{E}(\mathfrak{n},R)[U_{\mathfrak{p}}]$; cf. Eq. (7) and Lem. 6.3. Utilizing Thm. 6.5, we demonstrate that $E_{\mathfrak{n}}$ generates $\mathcal{E}(\mathfrak{n},R)[U_{\mathfrak{p}}]$ over $R$. Considering the values of $E_{\mathfrak{n}}$ on specific edges of $\mathscr{T}$, we observe that the subgroup $\mathcal{E}_{0}(\mathfrak{n},R)[U_{\mathfrak{p}}]$ of $\mathcal{E}(\mathfrak{n},R)[U_{\mathfrak{p}}]$, consisting of cuspidal harmonic cochains, is contained in $R\left[M(\mathfrak{p})\right]\cdot E_{\mathfrak{n}}$, where $R\left[M(\mathfrak{p})\right]:=\{c\in R\mid M(\mathfrak{p})\cdot c=0\}$. By [12, Lem. 7.2], let $R:=\mathbb{Z}/\ell^{k}\mathbb{Z}$ for a prime $\ell\nmid q(q-1)$ and a sufficiently large $k$, then there exists an injective map:
$\mathcal{C}(\mathfrak{p}^{r})_{\ell}[U_{\mathfrak{p}}]\subseteq\mathcal{T}(\mathfrak{p}^{r})_{\ell}[U_{\mathfrak{p}}]\hookrightarrow\mathcal{E}_{0}(\mathfrak{p}^{r},R)[U_{\mathfrak{p}}]\subseteq R\left[M(\mathfrak{p})\right]\cdot E_{\mathfrak{n}}\cong\frac{\mathbb{Z}_{\ell}}{M(\mathfrak{p})\mathbb{Z}_{\ell}}.$ (1)
Moreover, according to Thm. 1.4 and Sec. 5, we construct explicitly a cyclic subgroup of $\mathcal{C}(\mathfrak{n})_{\ell}[U_{\mathfrak{p}}]$ that is isomorphic to $\mathbb{Z}_{\ell}/M(\mathfrak{p})\mathbb{Z}_{\ell}$; see Eq. (4), Sequence (6), and Prop. 6.9. Hence, we demonstrate that $\mathcal{C}(\mathfrak{p}^{r})_{\ell}[U_{\mathfrak{p}}]=\mathcal{T}(\mathfrak{p}^{r})_{\ell}[U_{\mathfrak{p}}]$.
Remark 1.9.
The idea of considering Eq. (3) and Sequence (1) comes from the proof of $\mathcal{C}(\mathfrak{p})=\mathcal{T}(\mathfrak{p})$ for the prime level case in [11, Thm. 1.4]. More precisely, Pál showed that the image of the $\ell$-primary part of $\mathcal{T}(\mathfrak{p})$, for a prime $\ell\neq p$, in the group $\Phi_{\infty}$ of connected components of the Néron model of $J_{0}(\mathfrak{p})$ at $\infty$ with respect to specialization, injects into $\mathcal{E}_{0}(\mathfrak{p},\mathbb{Z}/\ell^{k}\mathbb{Z})$ for large enough $k$; cf. [11, p. 133], which leads to [11, Thm. 7.19].
Remark 1.10.
The primes $\ell$ dividing $q(q-1)$ lie beyond the scope of our current approach. Considering a non-zero ideal $\mathfrak{n}\lhd A$, $J_{0}(\mathfrak{n})$ has bad reduction both at $\infty$ and at the primes dividing $\mathfrak{n}$. In particular, $J_{0}(\mathfrak{n})$ has split toric reduction at $\infty$. In instances where $\ell$ divides $q(q-1)$, the specialization map $\mathcal{T}(\mathfrak{n})_{\ell}\rightarrow(\Phi_{\infty})_{\ell}$ might not ensure injectivity; cf. [12, Lem. 7.2].
The paper is outlined in the following. We begin in Sec. 2 with an introduction to the rational cuspidal divisor class group of $X_{0}(\mathfrak{n})$. Sec. 3 covers harmonic cochains on the Bruhat-Tits tree $\mathscr{T}$ of $\operatorname{PGL}(2,K_{\infty})$ and their Fourier expansions. Following this, Sec. 4 discusses Hecke operators acting on the space of $R$-valued harmonic cochains on $\mathscr{T}$, where $R$ is a commutative ring with unity, alongside Eisenstein ideals for prime-power levels. Sec. 5 explores the action of Hecke operators on the Jacobian $J_{0}(\mathfrak{n})$ of the Drifeld modular curve $X_{0}(\mathfrak{n})$. Finally, all introduced tools are integrated into the proof of Main Theorem.
2 The rational cuspidal divisor class group of $X_{0}(\mathfrak{n})$
In this section, we recall from [8]. Let $\mathfrak{n}\lhd A$ be a non-zero ideal. We also denote by $\mathfrak{n}$ the monic polynomial in $A$ generating the ideal. The rational cuspidal divisor class group $\mathcal{C}(\mathfrak{n})$ of $X_{0}(\mathfrak{n})$ is a subgroup of $J_{0}(\mathfrak{n})$ generated by the linear equivalence classes of the degree 00 rational cuspidal divisors on $X_{0}(\mathfrak{n})$. By Gekeler [5, (3.6)], the cusps of $X_{0}(\mathfrak{n})$ are in bijection with $\Gamma_{0}(\mathfrak{n})\backslash\mathbb{P}^{1}(K)$. Moreover, every cusp of $X_{0}(\mathfrak{n})$ has a representative $\begin{bmatrix}a\\ d\end{bmatrix}$ in $\Gamma_{0}(\mathfrak{n})\backslash\mathbb{P}^{1}(K)$, where $a,d\in A$ are monic, $d|\mathfrak{n}$, and $\gcd(a,\mathfrak{n})=1$. Such a cusp is called of height $d$, independent to the choice of $a$. Denote by $[0]:=\begin{bmatrix}1\\ 1\end{bmatrix}$ and $[\infty]:=\begin{bmatrix}1\\ \mathfrak{n}\end{bmatrix}$ the unique cusps of $X_{0}(\mathfrak{n})$ of height $1$ and $\mathfrak{n}$, respectively.
In the following, we fix a prime-power ideal $\mathfrak{n}=\mathfrak{p}^{r}\lhd A$ with $r\geq 1$. The cusps of $X_{0}(\mathfrak{n})$ of the same height are conjugate over $K$ by [13, Lem. 3.1]. For a monic $d\mid\mathfrak{n}$, let $(P(\mathfrak{n})_{d})$ be the sum of all the cusps of $X_{0}(\mathfrak{n})$ of height $d$, which is a rational cuspidal divisor on $X_{0}(\mathfrak{n})$; cf. [8, Sec. 1.3]. Let
$\operatorname{Div}_{\operatorname{cusp}}^{0}(X_{0}(\mathfrak{n}))(K):=\left\{C=\sum_{\begin{subarray}{c}d|\mathfrak{n}\\ \text{monic}\end{subarray}}a_{d}\cdot(P(\mathfrak{n})_{d})\bigm{|}\deg(C)=0,~{}a_{d}\in\mathbb{Z}\right\}$
be the group of the degree 00 rational cuspidal divisors on $X_{0}(\mathfrak{n})$. Recall that a modular function on $X_{0}(\mathfrak{n})$ is a meromorphic function on $\Omega\cup\ \mathbb{P}^{1}(K)$ which is invariant under $\Gamma_{0}(\mathfrak{n})$, and a modular unit on $X_{0}(\mathfrak{n})$ is a modular function on $X_{0}(\mathfrak{n})$ that does not have zeros or poles on $\Omega$. Let $\mathcal{U}_{\mathfrak{n}}$ be the subgroup of $\operatorname{Div}_{\operatorname{cusp}}^{0}(X_{0}(\mathfrak{n}))(K)$ consisting of the divisors of modular units. Let $C_{i}:=(P(\mathfrak{n})_{\mathfrak{p}^{i}})-\deg(P(\mathfrak{n})_{\mathfrak{p}^{i}})\cdot[\infty]\in\operatorname{Div}_{\operatorname{cusp}}^{0}(X_{0}(\mathfrak{n}))(K)$, where $10\leq i\leq r-$ 1. Then
$\mathcal{C}(\mathfrak{n}):=\operatorname{Div}_{\operatorname{cusp}}^{0}(X_{0}(\mathfrak{n}))(K)/\mathcal{U}_{\mathfrak{n}}=\langle\overline{C_{0}},\overline{C_{1}},\cdots,\overline{C_{r-1}}\rangle.$
Let $\Delta(z)$ be the Drinfeld discriminant function defined in [5, p. 183]. For $d|\mathfrak{n}$, consider $\Delta_{d}(z):=\Delta(dz)$, which is a modular form on $\Omega$ of weight $q^{2}-1$ and type 00 for $\Gamma_{0}(\mathfrak{n})$; cf. [5, (1.2)]. The zero orders of $\Delta_{d}(z)$ at the cusps of $X_{0}(\mathfrak{n})$ are defined in [2, p. 47]. Using [5, Eq. (3.10) and (3.11)], we found the divisor of $\Delta_{d}(z)$ on $X_{0}(\mathfrak{n})$ in [8, Sec. 2.1]. Let $\mathcal{E}_{\mathfrak{n}}$ be the group of modular units on $X_{0}(\mathfrak{n})$. Recall the following group homomorphism in [8, Sec. 2.1]:
${g:}$${\operatorname{Div}_{\operatorname{cusp}}^{0}(X_{0}(\mathfrak{n}))(K)}$$\tens{\displaystyle\mathcal{E}_{\mathfrak{n}}\tens_{\mathbb{Z}}\mathbb{Q}}$${C=\displaystyle\sum_{\begin{subarray}{c}d|\mathfrak{n}\\ \text{monic}\end{subarray}}a_{d}\cdot(P(\mathfrak{n})_{d})}$${\left(\displaystyle\prod_{\begin{subarray}{c}d|\mathfrak{n}\\ \text{monic}\end{subarray}}\Delta_{d}^{r_{d}}\right)\otimes\frac{1}{(q-1)(|\mathfrak{p}|^{2}-1)|\mathfrak{p}|^{r-1}},}$
where $r_{d}\in\mathbb{Z}$ are defined by
$\begin{bmatrix}r_{1}\\ r_{\mathfrak{p}}\\ \vdots\\ r_{\mathfrak{p}^{r}}\end{bmatrix}:=(q-1)(|\mathfrak{p}|^{2}-1)|\mathfrak{p}|^{r-1}\cdot\Lambda(\mathfrak{n})^{-1}\cdot\begin{bmatrix}a_{1}\\ a_{\mathfrak{p}}\\ \vdots\\ a_{\mathfrak{p}^{r}}\end{bmatrix}.$
Here, $\Lambda(\mathfrak{n})^{-1}\in\operatorname{GL}_{r+1}(\mathbb{Q})$ is a certain matrix defined in [8, Sec. 2.1]. The map $g$ provides a connection between the rational cuspidal divisor class group $\mathcal{C}(\mathfrak{n})$ and the modular units on $X_{0}(\mathfrak{n})$. More precisely, if $g(C)=f\otimes a$ for some $C\in\operatorname{Div}_{\operatorname{cusp}}^{0}(X_{0}(\mathfrak{n}))(K)$, $f\in\mathcal{E}_{\mathfrak{n}}$, and $a\in\mathbb{Q}$, then we have $C=a\cdot\operatorname{div}(f)$.
3 Harmonic cochains and Fourier expansions
Let $\mathscr{T}$ be the Bruhat-Tits tree of $\operatorname{PGL}(2,K_{\infty})$; cf. [7, (1.3)]. Recall that $G$ is the group scheme $\operatorname{GL}(2)$ over $\mathbb{F}_{q}$, with $Z$ the group scheme of scalar matrices in $G$. Let $\mathcal{K}=G(\mathcal{O}_{\infty})$ with its Iwahori subgroup $\mathcal{I}$ defined by
=$\mathcal{I}=\left\{\begin{pmatrix}a&b\\ c&d\end{pmatrix}\in\mathcal{K}\bigm{|}c\equiv 0\bmod\pi_{\infty}\right\}.$
Let $V(\mathscr{T})=G(K_{\infty})/\mathcal{K}\cdot Z(K_{\infty})$ and $E(\mathscr{T})=G(K_{\infty})/\mathcal{I}\cdot Z(K_{\infty})$ be the sets of the vertices and the oriented edges of $\mathscr{T}$, respectively; cf. [7, (1.3)]. The group $G(K_{\infty})$ acts on $E(\mathscr{T})$ by left multiplication. For a function $f$ on $E(\mathscr{T})$ and $\gamma\in G(K_{\infty})$, we define a function $f|\gamma$ on $E(\mathscr{T})$ by $(f|\gamma)(e):=f(\gamma e)$.
Definition 3.1.
(van der Put) Let $R$ be a commutative ring with unity. An $R$-valued harmonic cochain on $\mathscr{T}$ is a function $f:E(\mathscr{T})\rightarrow R$ that satisfies
  1. 1.
    $f(e)+f(\overline{e})=0\text{ for all }e\in E(\mathscr{T}).$
  2. 2.
    $\sum_{\begin{subarray}{c}e\in E(\mathscr{T})\\ t(e)=v\end{subarray}}f(e)=0\text{ for all }v\in V(\mathscr{T}).$
Here, for $e\in E(\mathscr{T})$, $t(e)$ is its terminus and $\overline{e}$ is its inversely oriented edge.
We introduce some notation. Fix a non-zero ideal $\mathfrak{n}\lhd A$. Denote by $\mathcal{H}(\mathscr{T},R)$ the group of $R$-valued harmonic cochains on $\mathscr{T}$ and by $\mathcal{H}(\mathfrak{n},R):=\mathcal{H}(\mathscr{T},R)^{\Gamma_{0}(\mathfrak{n})}$ its subgroup of $\Gamma_{0}(\mathfrak{n})$-invariant harmonic cochains, i.e. $f|\gamma=f$ for all $\gamma\in\Gamma_{0}(\mathfrak{n})$. For $f\in\mathcal{H}(\mathfrak{n},R)$, it naturally defines a function $f^{\prime}$ on the quotient graph $\Gamma_{0}(\mathfrak{n})\backslash\mathscr{T}$. If $f^{\prime}$ has compact support, then we call $f$ a cuspidal harmonic cochain for $\Gamma_{0}(\mathfrak{n})$. Let $\mathcal{H}_{0}(\mathfrak{n},R)$ be the subgroup of $\mathcal{H}(\mathfrak{n},R)$ consisting of cuspidal harmonic cochains for $\Gamma_{0}(\mathfrak{n})$, and denote by $\mathcal{H}_{00}(\mathfrak{n},R)$ the image of $\mathcal{H}_{0}(\mathfrak{n},\mathbb{Z})\otimes R$ in $\mathcal{H}_{0}(\mathfrak{n},R)$. Note that if $R$ is flat over $\mathbb{Z}$, then $\mathcal{H}_{0}(\mathfrak{n},R)=\mathcal{H}_{00}(\mathfrak{n},R)$.
Definition 3.2 (Pál [11, Sec. 2]).
A ring $R$ is called a coefficient ring if $1/p\in R$ and $R$ is a quotient of a discrete valuation ring $\widetilde{R}$ which contains $p$-th roots of unity.
In the following, we fix $R=\mathbb{C}$ or a coefficient ring. Let $f\in\mathcal{H}(\mathscr{T},R)$ which is invariant under the action of $\Gamma_{\infty}:=\left\{\begin{pmatrix}a&b\\ 0&d\end{pmatrix}\in G(A)\right\}$. The constant Fourier coefficient of $f$ is the $R$-valued function $f^{0}$ on $\pi_{\infty}^{\mathbb{Z}}$ defined by
$f^{0}(\pi_{\infty}^{k})=\begin{cases}q^{1-k}\sum_{u\in(\pi_{\infty})/(\pi_{\infty}^{k})}f\left(\begin{pmatrix}\pi_{\infty}^{k}&u\\ 0&1\end{pmatrix}\right)&\text{if $k\geq 1$};\\ f\left(\begin{pmatrix}\pi_{\infty}^{k}&0\\ 0&1\end{pmatrix}\right)&\text{if $k\leq 1$}.\end{cases}$
Let $\eta:K_{\infty}\rightarrow\mathbb{C}^{\times}$ be the character
$\eta:\sum a_{i}\pi_{\infty}^{i}\mapsto\eta_{0}(\operatorname{Tr}(a_{1})),$
where $\operatorname{Tr}:\mathbb{F}_{q}\rightarrow\mathbb{F}_{p}$ is the trace and $\eta_{0}$ is any non-trivial character of $\mathbb{F}_{p}$. For a divisor $\mathfrak{m}$ on $K$, the $\mathfrak{m}$-th Fourier coefficient $f^{\ast}(\mathfrak{m})$ of $f$ is
$f^{\ast}(\mathfrak{m})=q^{-1-\deg(\mathfrak{m})}\sum_{u\in(\pi_{\infty})/(\pi_{\infty}^{2+\deg(\mathfrak{m})})}f\left(\begin{pmatrix}\pi_{\infty}^{2+\deg(\mathfrak{m})}&u\\ 0&1\end{pmatrix}\right)\eta(-mu),$
if $\mathfrak{m}$ is non-negative; $f^{\ast}(\mathfrak{m})=0$, otherwise. Here, $m\in A$ is the monic polynomial such that $\mathfrak{m}=\operatorname{div}(m)\cdot\infty^{\deg(\mathfrak{m})}$.
Theorem 3.3 (Gekeler [4, (2.8)]).
Let $f\in\mathcal{H}(\mathscr{T},R)$ which is invariant under $\Gamma_{\infty}$. Then
$f\left(\begin{pmatrix}\pi_{\infty}^{k}&y\\ 0&1\end{pmatrix}\right)=f^{0}(\pi_{\infty}^{k})+\sum_{\begin{subarray}{c}0\neq m\in A\\ \deg(m)\leq k-2\end{subarray}}f^{\ast}(\operatorname{div}(m)\cdot\infty^{k-2})\cdot\eta(my).$
In particular, $f$ is uniquely determined by the Fourier coefficients $f^{0}$ and $f^{\ast}$.
Recall the following theorem:
Theorem 3.4 (van der Put [15, Prop. (1.1)]).
Let $\mathcal{O}(\Omega)^{\ast}$ be the group of non-vanishing holomorphic rigid-analytic functions on $\Omega$. There is a canonical exact sequence of $G(K_{\infty})$-modules
$0\rightarrow\mathbb{C}_{\infty}^{\ast}\rightarrow\mathcal{O}(\Omega)^{\ast}\overset{\widetilde{r}}{\rightarrow}\mathcal{H}(\mathscr{T},\mathbb{Z})\rightarrow 0$
The van der Put map $\widetilde{r}$ extends naturally to the map
[…]
Example.
Recall that $\Delta\in\mathcal{O}(\Omega)^{\ast}$ is the Drinfeld discriminant function. Consider the function $\widetilde{r}(\Delta)\in\mathcal{H}(\mathscr{T},\mathbb{Z})\subset\mathcal{H}(\mathscr{T},\mathbb{C})$, which is invariant under $\Gamma_{\infty}$. By Gekeler [5, (2.5) and Cor. 2.8], we have
$\begin{cases}\widetilde{r}(\Delta)^{0}(\pi_{\infty}^{k})=-(q-1)q^{1-k};\\ \widetilde{r}(\Delta)^{\ast}(1)=(q^{2}-1)(q-1)/q.\end{cases}$ (2)
Recall the following lemma:
Lemma 3.5 ([14, Lem. 2.14]).
Given a non-zero ideal $\mathfrak{m}\lhd A$ and $f\in\mathcal{H}(\mathscr{T},R)$ which is invariant under $\Gamma_{\infty}$, define
$f|B_{\mathfrak{m}}:=f|\begin{pmatrix}\mathfrak{m}&0\\ 0&1\end{pmatrix}\text{~{}with~{}}f|B_{\mathfrak{m}}^{-1}:=f|\begin{pmatrix}1&0\\ 0&\mathfrak{m}\end{pmatrix}.$
Then we have
$\begin{cases}(f|B_{\mathfrak{m}})^{0}(\pi_{\infty}^{k})=f^{0}(\pi_{\infty}^{k-\deg(\mathfrak{m})});\\ (f|B_{\mathfrak{m}})^{\ast}(\mathfrak{n})=f^{\ast}(\mathfrak{n}/\mathfrak{m}).\end{cases}$
Then we derive that:
Lemma 3.6.
Fix a prime ideal $\mathfrak{p}\lhd A$ and $i\geq 1$. Consider $\widetilde{r}(\Delta_{\mathfrak{p}^{i}})\in\mathcal{H}(\mathscr{T},\mathbb{Z})\subset\mathcal{H}(\mathscr{T},\mathbb{C})$, which is invariant under $\Gamma_{\infty}$. Then we have
$\begin{cases}\widetilde{r}(\Delta_{\mathfrak{p}^{i}})^{0}(\pi_{\infty}^{k})=-(q-1)q^{1-k}|\mathfrak{p}|^{i};\\ \widetilde{r}(\Delta_{\mathfrak{p}^{i}})^{\ast}(1)=0.\end{cases}$
Proof..
By the previous lemma, we have $\widetilde{r}(\Delta_{\mathfrak{p}^{i}})^{0}(\pi_{\infty}^{k})=\widetilde{r}(\Delta)^{0}(\pi_{\infty}^{k-\deg(\mathfrak{p}^{i})})$ and $\widetilde{r}(\Delta_{\mathfrak{p}^{i}})^{\ast}(1)=\widetilde{r}(\Delta)^{\ast}((\mathfrak{p}^{i})^{-1})$, which imply the result by Eq. (2). ∎
4 Hecke operators and Eisenstein ideals
Fix a prime-power ideal $\mathfrak{n}=\mathfrak{p}^{r}$ of $A$ with $r\geq 2$, and let $R$ be a commutative ring with unity. For a non-zero ideal $\mathfrak{m}\lhd A$, we define an $R$-linear transformation of the space of $R$-valued functions on $E(\mathscr{T})$ by
$f|T_{\mathfrak{m}}:=\sum f|\begin{pmatrix}a&b\\ 0&d\end{pmatrix},$
where the sum is over $a,b,d\in A$ such that $a$, $d$ are monic, $(ad)=\mathfrak{m}$, $(a)+\mathfrak{n}=A$, and $\deg(b)<\deg(d)$. This transformation is the $\mathfrak{m}$-th Hecke operator; cf. [13, Sec. 2.3]. In particular, for the prime divisor $\mathfrak{p}$ of $\mathfrak{n}$, write
$f|U_{\mathfrak{p}}:=\sum_{\begin{subarray}{c}b\in A\\ \deg(b)<\deg(\mathfrak{p})\end{subarray}}f|\begin{pmatrix}1&b\\ 0&\mathfrak{p}\end{pmatrix},$
then we see that $U_{\mathfrak{p}}=T_{\mathfrak{p}}$.
Remark 4.1.
The definition of $T_{\mathfrak{p}}:=U_{\mathfrak{p}}$ here is different from that of $T_{\mathfrak{p}}:=U_{\mathfrak{p}}+B_{\mathfrak{p}}$ in [5, (1.17)].
Proposition 4.2 ([12, Prop. 2.10]).
The Hecke operators preserve the spaces $\mathcal{H}(\mathfrak{n},R)$ and $\mathcal{H}_{0}(\mathfrak{n},R)$, and satisfy the recursive formulas:
  1. 1.
    $T_{\mathfrak{m}\mathfrak{m}^{\prime}}=T_{\mathfrak{m}}T_{\mathfrak{m}^{\prime}}$ if $\mathfrak{m}+\mathfrak{m}^{\prime}=A$.
  2. 2.
    $T_{\mathfrak{q}^{i}}=T_{\mathfrak{q}^{i-1}}T_{\mathfrak{q}}-|\mathfrak{q}|T_{\mathfrak{q}^{i-2}}$ for any prime $\mathfrak{q}\neq\mathfrak{p}$ and $i\geq 2$.
  3. 3.
    $T_{\mathfrak{p}^{i}}=U_{\mathfrak{p}}^{i}$ for $i\geq 1$.
Example.
Consider the action of $U_{\mathfrak{p}}$ on $\mathcal{H}(\mathfrak{n},R)$. Let $f\in\mathcal{H}(\mathfrak{n},R)$. Since $\mathfrak{p}^{2}\mid\mathfrak{n}$, we have $h:=f|U_{\mathfrak{p}}\in\mathcal{H}(\mathfrak{n}/\mathfrak{p},R)$ by [13, Lem. 2.6]. Moreover, since $h$ is invariant under $\Gamma_{\infty}$, we have $(h|B_{\mathfrak{p}})|U_{\mathfrak{p}}=|\mathfrak{p}|\cdot h$ by [12, Lem. 2.23]. We obtain:
${U_{\mathfrak{p}}:\mathcal{H}(\mathfrak{n},R)}$${\mathcal{H}(\mathfrak{n}/\mathfrak{p},R)\subset\mathcal{H}(\mathfrak{n},R)}$${f}$${h:=f|U_{\mathfrak{p}}}$${h|B_{\mathfrak{p}}}$${{|\mathfrak{p}|\cdot h}.}$
Lemma 4.3 ([5, Cor. 2.11]).
Let $\mathfrak{q}\lhd A$ be a prime not equal to $\mathfrak{p}$. Then
  1. 1.
    $\widetilde{r}(\Delta)|T_{\mathfrak{q}}=(|\mathfrak{q}|+1)\cdot\widetilde{r}(\Delta).$
  2. 2.
    $\widetilde{r}(\Delta)|(U_{\mathfrak{p}}+B_{\mathfrak{p}})=(|\mathfrak{p}|+1)\cdot\widetilde{r}(\Delta).$ Alternatively,
    $\widetilde{r}(\Delta)|U_{\mathfrak{p}}=(|\mathfrak{p}|+1)\cdot\widetilde{r}(\Delta)-\widetilde{r}(\Delta_{\mathfrak{p}}).$
From the above, we prove the following:
Lemma 4.4.
Let $k\geq 1$ and $\mathfrak{q}\lhd A$ be a prime not equal to $\mathfrak{p}$. Then
  1. 1.
    $\widetilde{r}(\Delta_{\mathfrak{p}^{k}})|T_{\mathfrak{q}}=(|\mathfrak{q}|+1)\cdot\widetilde{r}(\Delta_{\mathfrak{p}^{k}})$.
  2. 2.
    $\widetilde{r}(\Delta_{\mathfrak{p}^{k}})|U_{\mathfrak{p}}=|\mathfrak{p}|\cdot\widetilde{r}(\Delta_{\mathfrak{p}^{k-1}})$.
Proof..
First, we have
$\widetilde{r}(\Delta_{\mathfrak{p}^{k}})|T_{\mathfrak{q}}=\widetilde{r}(\Delta)|\begin{pmatrix}\mathfrak{p}^{k}&0\\ 0&1\end{pmatrix}\begin{pmatrix}\mathfrak{q}&0\\ 0&1\end{pmatrix}+\sum_{\begin{subarray}{c}b\in A\\ \deg(b)<\deg(\mathfrak{q})\end{subarray}}\widetilde{r}(\Delta)|\begin{pmatrix}\mathfrak{p}^{k}&0\\ 0&1\end{pmatrix}\begin{pmatrix}1&b\\ 0&\mathfrak{q}\end{pmatrix}$
$=\widetilde{r}(\Delta)|\begin{pmatrix}\mathfrak{q}&0\\ 0&1\end{pmatrix}\begin{pmatrix}\mathfrak{p}^{k}&0\\ 0&1\end{pmatrix}+\sum_{\begin{subarray}{c}b\in A\\ \deg(b)<\deg(\mathfrak{q})\end{subarray}}\widetilde{r}(\Delta)|\begin{pmatrix}1&b\mathfrak{p}^{k}\\ 0&\mathfrak{q}\end{pmatrix}\begin{pmatrix}\mathfrak{p}^{k}&0\\ 0&1\end{pmatrix}$
$=\widetilde{r}(\Delta)|\begin{pmatrix}\mathfrak{q}&0\\ 0&1\end{pmatrix}\begin{pmatrix}\mathfrak{p}^{k}&0\\ 0&1\end{pmatrix}+\sum_{\begin{subarray}{c}b\in A\\ \deg(b)<\deg(\mathfrak{q})\end{subarray}}\widetilde{r}(\Delta)|\begin{pmatrix}1&b\\ 0&\mathfrak{q}\end{pmatrix}\begin{pmatrix}\mathfrak{p}^{k}&0\\ 0&1\end{pmatrix}$
$=(\widetilde{r}(\Delta)|T_{\mathfrak{q}})|\begin{pmatrix}\mathfrak{p}^{k}&0\\ 0&1\end{pmatrix}=(|\mathfrak{q}|+1)\cdot\widetilde{r}(\Delta)|\begin{pmatrix}\mathfrak{p}^{k}&0\\ 0&1\end{pmatrix}=(|\mathfrak{q}|+1)\cdot\widetilde{r}(\Delta_{\mathfrak{p}^{k}}).$
For the second statement, since $\widetilde{r}(\Delta_{\mathfrak{p}^{k-1}})$ is $\Gamma_{\infty}$-invariant, we have
$\widetilde{r}(\Delta_{\mathfrak{p}^{k}})|U_{\mathfrak{p}}=\sum_{\begin{subarray}{c}b\in A\\ \deg(b)<\deg(\mathfrak{p})\end{subarray}}\widetilde{r}(\Delta_{\mathfrak{p}^{k-1}})|\begin{pmatrix}\mathfrak{p}&0\\ 0&1\end{pmatrix}\begin{pmatrix}1&b\\ 0&\mathfrak{p}\end{pmatrix}$
$=\sum_{\begin{subarray}{c}b\in A\\ \deg(b)<\deg(\mathfrak{p})\end{subarray}}\widetilde{r}(\Delta_{\mathfrak{p}^{k-1}})|\begin{pmatrix}\mathfrak{p}&0\\ 0&\mathfrak{p}\end{pmatrix}\begin{pmatrix}1&b\\ 0&1\end{pmatrix}=|\mathfrak{p}|\cdot\widetilde{r}(\Delta_{\mathfrak{p}^{k-1}}).$
Recall that $\mathbb{T}(\mathfrak{n})$ is the $\mathbb{Z}$-algebra generated by all Hecke operators $T_{\mathfrak{q}}$, where $\mathfrak{q}\lhd A$ are primes, acting on the group $\mathcal{H}_{0}(\mathfrak{n},\mathbb{Z})$, and the Eisenstein ideal $\mathfrak{E}(\mathfrak{n})$ of $\mathbb{T}(\mathfrak{n})$ is the ideal generated by the elements
$\{T_{\mathfrak{q}}-|\mathfrak{q}|-1\mid\mathfrak{q}\text{ is prime},\mathfrak{q}\nmid\mathfrak{n}\}.$
To simplify the notation, we put
  • $\mathcal{E}(\mathfrak{n},R):=\mathcal{H}(\mathfrak{n},R)[\mathfrak{E}(\mathfrak{n})]\subset\mathcal{H}(\mathfrak{n},R)$,
  • $\mathcal{E}_{0}(\mathfrak{n},R):=\mathcal{H}_{0}(\mathfrak{n},R)[\mathfrak{E}(\mathfrak{n})]\subset\mathcal{H}_{0}(\mathfrak{n},R)$,
  • $\mathcal{E}_{00}(\mathfrak{n},R):=\mathcal{H}_{00}(\mathfrak{n},R)[\mathfrak{E}(\mathfrak{n})]\subset\mathcal{H}_{00}(\mathfrak{n},R)$,
the subgroups consisting of the elements annihilated by $\mathfrak{E}(\mathfrak{n})$.
Let $v$ be a place of $K$. Denote by $K_{v}$ the completion of $K$ at $v$, $\mathcal{O}_{v}$ the ring of integers of $K_{v}$, and $\mathbb{F}_{v}$ the residue field of $\mathcal{O}_{v}$. Moreover, we denote by $\mathbb{F}_{v}^{\text{nr}}$ and $\mathcal{O}_{v}^{\text{nr}}$ the maximal unramified extension of $\mathbb{F}_{v}$ and $\mathcal{O}_{v}$, respectively. Let $\mathcal{J}$ denote the Néron model of $J_{0}(\mathfrak{n})$ over $\mathcal{O}_{v}$. Let $\mathcal{J}^{0}$ denote the relative connected component of the identity of $\mathcal{J}$, that is, the largest open subscheme of $\mathcal{J}$ containing the identity section which has connected fibres. The group of connected components of $J_{0}(\mathfrak{n})$ at $v$ is $\Phi_{v}:=\mathcal{J}_{\mathbb{F}_{v}}/\mathcal{J}_{\mathbb{F}_{v}}^{0}$. There is a natural map in [12, p. 586], called the canonical specialization map:
$\rho_{v}:J_{0}(\mathfrak{n})(K_{v}^{\text{nr}})=\mathcal{J}(\mathcal{O}_{v}^{\text{nr}})\rightarrow\mathcal{J}_{\mathbb{F}_{v}}(\overline{\mathbb{F}_{v}})\rightarrow\Phi_{v}.$ (3)
The Hecke algebra $\mathbb{T}(\mathfrak{n})$ naturally acts on the Jacobian variety $J_{0}(\mathfrak{n})$ of $X_{0}(\mathfrak{n})$, which functorially extends to $\mathcal{J}$; hence, $\mathbb{T}(\mathfrak{n})$ also acts on $\Phi_{v}$; cf. [13, p. 386].
5 Two degeneracy maps
Fix a prime-power ideal $\mathfrak{n}=\mathfrak{p}^{r}$ of $A$ with $r\geq 2$. Recall that the curve $Y_{0}(\mathfrak{n})$ is the generic fibre of the coarse moduli scheme for the functor which associates to an $A$-scheme $S$ the set of isomorphism classes of pairs $(\phi,C_{\mathfrak{n}})$, where $\phi$ is a rank-$2$ Drinfeld module over $S$ with an $\mathfrak{n}$-cyclic subgroup $C_{\mathfrak{n}}$; for details, see [1], [3], and [7]. Let $\alpha_{\mathfrak{p}}(\mathfrak{n}),\beta_{\mathfrak{p}}(\mathfrak{n}):Y_{0}(\mathfrak{n})\rightarrow Y_{0}(\mathfrak{n}/\mathfrak{p})$ be the degeneracy maps with moduli-theoretic interpretations:
$\alpha:=\alpha_{\mathfrak{p}}(\mathfrak{n}):(\phi,C_{\mathfrak{n}})\mapsto(\phi,C_{\mathfrak{n}/\mathfrak{p}}),$
$\beta:=\beta_{\mathfrak{p}}(\mathfrak{n}):(\phi,C_{\mathfrak{n}})\mapsto(\phi/C_{\mathfrak{p}},C_{\mathfrak{n}}/C_{\mathfrak{p}}),$
where $C_{\mathfrak{n}/\mathfrak{p}}$ and $C_{\mathfrak{p}}$ are the subgroups of $C_{\mathfrak{n}}$ of order $\mathfrak{n}/\mathfrak{p}$ and $\mathfrak{p}$, respectively. These morphisms are proper, and hence uniquely extend to morphisms $\alpha,\beta:X_{0}(\mathfrak{n})\rightarrow X_{0}(\mathfrak{n}/\mathfrak{p})$; cf. [12, p. 584]. Moreover, by Picard functoriality, the maps $\alpha$ and $\beta$ induce two homomorphisms:
$\alpha^{\ast}:=\alpha_{\mathfrak{p}}(\mathfrak{n})^{\ast}:J_{0}(\mathfrak{n}/\mathfrak{p})\rightarrow J_{0}(\mathfrak{n}),~{}\beta^{\ast}:=\beta_{\mathfrak{p}}(\mathfrak{n})^{\ast}:J_{0}(\mathfrak{n}/\mathfrak{p})\rightarrow J_{0}(\mathfrak{n}),$
with their duals denoted by
$\alpha_{\ast}:=\alpha_{\mathfrak{p}}(\mathfrak{n})_{\ast}:J_{0}(\mathfrak{n})\rightarrow J_{0}(\mathfrak{n}/\mathfrak{p}),~{}\beta_{\ast}:=\beta_{\mathfrak{p}}(\mathfrak{n})_{\ast}:J_{0}(\mathfrak{n})\rightarrow J_{0}(\mathfrak{n}/\mathfrak{p}).$
We take the Hecke operator “$U_{\mathfrak{p}}$” acting on $J_{0}(\mathfrak{n})$ by:
${U_{\mathfrak{p}}:J_{0}(\mathfrak{n})}$${J_{0}(\mathfrak{n}/\mathfrak{p})}$${J_{0}(\mathfrak{n}).}$$\scriptstyle{\beta_{\ast}}$$\scriptstyle{\alpha^{\ast}}$
Note that $U_{\mathfrak{p}}:=\alpha_{\mathfrak{p}}(\mathfrak{n})^{\ast}\circ\beta_{\mathfrak{p}}(\mathfrak{n})_{\ast}=\beta_{\mathfrak{p}}(\mathfrak{n}\mathfrak{p})_{\ast}\circ\alpha_{\mathfrak{p}}(\mathfrak{n}\mathfrak{p})^{\ast}$; cf. [17, Rem. 3.5]. The definition of $U_{\mathfrak{p}}$ here is the dual version of that in [12, p. 584], which are conjugate to each other by the Atkin-Lehner involution $W_{\mathfrak{n}}$; cf. [13, p. 398].
Remark 5.1.
For a prime ideal $\mathfrak{q}\lhd A$ not dividing $\mathfrak{n}$, the Hecke operator $T_{\mathfrak{q}}$ commutes with the Atkin–Lehner involution $W_{\mathfrak{n}}$, so $T_{\mathfrak{q}}$ coincides with its dual; cf. [12, p. 584] and [13, p. 398].
Denote $Y_{0}(\mathfrak{n})_{\mathbb{C}_{\infty}}:=\Gamma_{0}(\mathfrak{n})\backslash\Omega$ with its completion $X_{0}(\mathfrak{n})_{\mathbb{C}_{\infty}}:=\Gamma_{0}(\mathfrak{n})\backslash(\Omega\cup\mathbb{P}^{1}(K))$. For $z\in\Omega\cup\mathbb{P}^{1}(K)$, we have $\alpha,\beta:X_{0}(\mathfrak{n})_{\mathbb{C}_{\infty}}\rightarrow X_{0}(\mathfrak{n}/\mathfrak{p})_{\mathbb{C}_{\infty}}$ with
$\begin{cases}\alpha(\Gamma_{0}(\mathfrak{n})\cdot z)=\Gamma_{0}(\mathfrak{n}/\mathfrak{p})\cdot z;\\ \beta(\Gamma_{0}(\mathfrak{n})\cdot z)=\Gamma_{0}(\mathfrak{n}/\mathfrak{p})\cdot\begin{pmatrix}\mathfrak{p}&0\\ 0&1\end{pmatrix}\cdot z.\end{cases}$
Then for a cusp $\Gamma_{0}(\mathfrak{n})\cdot\begin{bmatrix}a\\ \mathfrak{p}^{i}\end{bmatrix}$ of $X_{0}(\mathfrak{n})_{\mathbb{C}_{\infty}}$ in the form defined in Sec. 2, we have
$\alpha(\Gamma_{0}(\mathfrak{n})\cdot\begin{bmatrix}a\\ \mathfrak{p}^{i}\end{bmatrix})=\begin{cases}\Gamma_{0}(\mathfrak{n}/\mathfrak{p})\cdot\begin{bmatrix}a\\ \mathfrak{p}^{i}\end{bmatrix},&\text{if $0\leq i\leq r-1$.}\\ \\ \Gamma_{0}(\mathfrak{n}/\mathfrak{p})\cdot\begin{bmatrix}1\\ \mathfrak{p}^{r-1}\end{bmatrix},&\text{if $i=r$.}\end{cases}$
$\beta(\Gamma_{0}(\mathfrak{n})\cdot\begin{bmatrix}a\\ \mathfrak{p}^{i}\end{bmatrix})=\begin{cases}\Gamma_{0}(\mathfrak{n}/\mathfrak{p})\cdot\begin{bmatrix}0\\ 1\end{bmatrix},&\text{if $i=0$.}\\ \\ \Gamma_{0}(\mathfrak{n}/\mathfrak{p})\cdot\begin{bmatrix}a\\ \mathfrak{p}^{i-1}\end{bmatrix},&\text{if $1\leq i\leq r$.}\end{cases}$
Moreover, the degeneracy coverings $\alpha$ and $\beta$ have degree $|\mathfrak{p}|$. From the above and [8, Lem. 2.2], $\alpha$ and $\beta$ also induce maps between the divisor groups:
$\alpha^{\ast},\beta^{\ast}:\operatorname{Div}(X_{0}(\mathfrak{n}/\mathfrak{p}))\rightarrow\operatorname{Div}(X_{0}(\mathfrak{n}));$
$\alpha_{\ast},\beta_{\ast}:\operatorname{Div}(X_{0}(\mathfrak{n}))\rightarrow\operatorname{Div}(X_{0}(\mathfrak{n}/\mathfrak{p})).$
Recall that $(P(\mathfrak{n})_{d})$ is the sum of all the cusps of $X_{0}(\mathfrak{n})$ of height $d$. We have:
Lemma 5.2.
  1. 1.
    For $(P(\mathfrak{n}/\mathfrak{p})_{\mathfrak{p}^{j}})\in\operatorname{Div}(X_{0}(\mathfrak{n}/\mathfrak{p}))$,
    $\alpha^{\ast}(P(\mathfrak{n}/\mathfrak{p})_{\mathfrak{p}^{j}})=\begin{cases}|\mathfrak{p}|\cdot(P(\mathfrak{n})_{\mathfrak{p}^{j}}),&\text{if $0\leq j\leq\lfloor\frac{r-1}{2}\rfloor$.}\\ (P(\mathfrak{n})_{\mathfrak{p}^{j}}),&\text{if $\lfloor\frac{r+1}{2}\rfloor\leq j\leq r-2$.}\\ (q-1)\cdot(P(\mathfrak{n})_{\mathfrak{p}^{r-1}})+(P(\mathfrak{n})_{\mathfrak{p}^{r}}),&\text{if $j=r-1$.}\end{cases}$
  2. 2.
    For $(P(\mathfrak{n})_{\mathfrak{p}^{j}})\in\operatorname{Div}(X_{0}(\mathfrak{n}))$,
    $\beta_{\ast}(P(\mathfrak{n})_{\mathfrak{p}^{j}})=\begin{cases}(P(\mathfrak{n}/\mathfrak{p})_{1}),&\text{if $j=0$.}\\ \frac{|\mathfrak{p}|-1}{q-1}\cdot(P(\mathfrak{n}/\mathfrak{p})_{1}),&\text{if $j=1$.}\\ |\mathfrak{p}|\cdot(P(\mathfrak{n}/\mathfrak{p})_{\mathfrak{p}^{j-1}}),&\text{if $2\leq j\leq\lfloor\frac{r}{2}\rfloor$.}\\ (P(\mathfrak{n}/\mathfrak{p})_{\mathfrak{p}^{j-1}}),&\text{if $\lfloor\frac{r}{2}\rfloor+1\leq j\leq r$.}\end{cases}$
Proof..
Inspired from Ex. Example, we decompose
${U_{\mathfrak{p}}:\mathcal{H}(\mathfrak{n},\mathbb{Z})}$${\mathcal{H}(\mathfrak{n}/\mathfrak{p},\mathbb{Z})}$${\mathcal{H}(\mathfrak{n},\mathbb{Z})}$${f}$${f|U_{\mathfrak{p}}}$${h}$${h.}$$\scriptstyle{\beta_{\ast}}$$\scriptstyle{\alpha^{\ast}}$
The first statement of lemma follows directly from [8, Lem. 2.1]. By Lem. 4.4, we have $\beta_{\ast}(\widetilde{r}(\Delta_{\mathfrak{p}})):=\widetilde{r}(\Delta_{\mathfrak{p}})|U_{\mathfrak{p}}=|\mathfrak{p}|\cdot\widetilde{r}(\Delta)$. Hence, we derive that $\beta_{\ast}$ maps $\operatorname{div}(\Delta_{\mathfrak{p}})\in\operatorname{Div}(X_{0}(\mathfrak{n}))$ to $|\mathfrak{p}|\cdot\operatorname{div}(\Delta)\in\operatorname{Div}(X_{0}(\mathfrak{n}/\mathfrak{p}))$. From [8, Sec. 2.1], the divisor of $\Delta_{\mathfrak{p}}$ on $X_{0}(\mathfrak{n})$ is
$\sum_{j=0}^{r}a_{j}\cdot(P(\mathfrak{n})_{\mathfrak{p}^{j}}):=|\mathfrak{p}|^{r-1}\cdot(P(\mathfrak{n})_{1})+\sum_{j=1}^{r}\frac{q-1}{\rho(j)}|\mathfrak{p}|^{\max\{r-2j+1,1\}}\cdot(P(\mathfrak{n})_{\mathfrak{p}^{j}}),$
where $a_{j}\in\mathbb{N}$ and
$\rho(j)=\begin{cases}1,&\text{if $0<j<r$.}\\ q-1,&\text{otherwise.}\end{cases}$
Additionally, the divisor of $\Delta^{|\mathfrak{p}|}$ on $X_{0}(\mathfrak{n}/\mathfrak{p})$ is
$\sum_{j=0}^{r-1}b_{j}\cdot(P(\mathfrak{n}/\mathfrak{p})_{\mathfrak{p}^{j}}):=|\mathfrak{p}|\cdot(P(\mathfrak{n}/\mathfrak{p})_{\mathfrak{p}^{r-1}})+\sum_{j=0}^{r-2}\frac{q-1}{\rho(j)}|\mathfrak{p}|^{\max\{r-2j,1\}}\cdot(P(\mathfrak{n}/\mathfrak{p})_{\mathfrak{p}^{j}}),$
where $b_{j}\in\mathbb{N}$. Now, the result for $2\leq j\leq r$ follows from $\beta_{\ast}(P(\mathfrak{n})_{\mathfrak{p}^{j}})=\frac{b_{j-1}}{a_{j}}\cdot(P(\mathfrak{n}/\mathfrak{p})_{\mathfrak{p}^{j-1}})$. For $j=0\text{ or }1$, we have $\beta_{\ast}(P(\mathfrak{n})_{1})=x\cdot(P(\mathfrak{n}/\mathfrak{p})_{1})$ and $\beta_{\ast}(P(\mathfrak{n})_{\mathfrak{p}})=y\cdot(P(\mathfrak{n}/\mathfrak{p})_{1})$ for some $x,y\in\mathbb{N}$. Consider
$f:=\left(\frac{\Delta}{\Delta_{\mathfrak{p}}}\right)^{|\mathfrak{p}|}\left(\frac{\Delta_{\mathfrak{p}^{2}}}{\Delta_{\mathfrak{p}}}\right)\in\mathcal{E}_{\mathfrak{n}}.$
Then $\widetilde{r}(f)\in\mathcal{H}(\mathfrak{n},\mathbb{Z})$. According to Lem. 4.3 and 4.4, we have $\widetilde{r}(f)|U_{\mathfrak{p}}=0$. Moreover, referencing [8, Sec. 2.1], the divisor of $f$ on $X_{0}(\mathfrak{n})$ is given by:
$(q-1)(|\mathfrak{p}|^{2}-1)|\mathfrak{p}|^{r-2}\cdot\left[\frac{|\mathfrak{p}|-1}{q-1}\cdot(P(\mathfrak{n})_{1})-(P(\mathfrak{n})_{\mathfrak{p}})\right].$
This implies that $y=\frac{|\mathfrak{p}|-1}{q-1}\cdot x$. Now, considering $a_{0}$, $a_{1}$, and $b_{0}$, we find
$|\mathfrak{p}|^{r-1}\cdot\beta_{\ast}(P(\mathfrak{n})_{1})+(q-1)|\mathfrak{p}|^{r-1}\cdot\beta_{\ast}(P(\mathfrak{n})_{\mathfrak{p}})=|\mathfrak{p}|^{r}\cdot(P(\mathfrak{n}/\mathfrak{p})_{1}),$
which implies that $x=1$ and $y=\frac{|\mathfrak{p}|-1}{q-1}$. ∎
In the following, we prove that $\mathcal{C}(\mathfrak{p}^{r-1})/\beta_{\ast}(\mathcal{C}(\mathfrak{p}^{r}))$ is $p$-primary.
Lemma 5.3 (cf. [17, Lem. 3.4]).
Fix a prime-power ideal $\mathfrak{n}=\mathfrak{p}^{r}\lhd A$ with $r\geq 2$. Let $\beta_{\ast}:\operatorname{Div}_{\operatorname{cusp}}^{0}(X_{0}(\mathfrak{p}^{r}))(K)\rightarrow\operatorname{Div}_{\operatorname{cusp}}^{0}(X_{0}(\mathfrak{p}^{r-1}))(K)$ be the map induced from Lem. 5.2. Then
$\operatorname{Div}_{\operatorname{cusp}}^{0}(X_{0}(\mathfrak{p}^{r-1}))(K)/\beta_{\ast}(\operatorname{Div}_{\operatorname{cusp}}^{0}(X_{0}(\mathfrak{p}^{r}))(K))\cong(\mathbb{Z}/|\mathfrak{p}|\mathbb{Z})^{\lfloor\frac{r}{2}\rfloor-1}.$
In particular, we have
$\mathcal{C}(\mathfrak{p}^{r-1})/\beta_{\ast}(\mathcal{C}(\mathfrak{p}^{r}))\hookrightarrow(\mathbb{Z}/|\mathfrak{p}|\mathbb{Z})^{\lfloor\frac{r}{2}\rfloor-1}.$
Proof..
Let $C_{i}:=(P(\mathfrak{n})_{\mathfrak{p}^{i}})-\deg(P(\mathfrak{n})_{\mathfrak{p}^{i}})\cdot[\infty]\in\operatorname{Div}_{\operatorname{cusp}}^{0}(X_{0}(\mathfrak{n}))(K)$ for $10\leq i\leq r-$ 1, and $C_{i}^{\prime}:=(P(\mathfrak{n}/\mathfrak{p})_{\mathfrak{p}^{i}})-\deg(P(\mathfrak{n}/\mathfrak{p})_{\mathfrak{p}^{i}})\cdot[\infty]\in\operatorname{Div}_{\operatorname{cusp}}^{0}(X_{0}(\mathfrak{n}/\mathfrak{p}))(K)$. Then $\operatorname{Div}_{\operatorname{cusp}}^{0}(X_{0}(\mathfrak{n}))(K)$ is generated by $\{C_{0},C_{1},\cdots,C_{r-1}\}$. By Lem. 5.2, we have
$\beta_{\ast}(\operatorname{Div}_{\operatorname{cusp}}^{0}(X_{0}(\mathfrak{n}))(K))$
$=\langle\beta_{\ast}(C_{0}),\beta_{\ast}(C_{1}),\beta_{\ast}(C_{2}),\cdots,\beta_{\ast}(C_{r-1})\rangle$
$=\langle C_{0}^{\prime},\frac{|\mathfrak{p}|-1}{q-1}C_{0}^{\prime},|\mathfrak{p}|C_{1}^{\prime},\cdots,|\mathfrak{p}|C_{\lfloor\frac{r}{2}\rfloor-1}^{\prime},C_{\lfloor\frac{r}{2}\rfloor}^{\prime},\cdots,C_{r-2}^{\prime}\rangle$
$\subset\operatorname{Div}_{\operatorname{cusp}}^{0}(X_{0}(\mathfrak{n}/\mathfrak{p}))(K)$
$=\langle C_{0}^{\prime},C_{1}^{\prime},\cdots,C_{\lfloor\frac{r}{2}\rfloor-1}^{\prime},C_{\lfloor\frac{r}{2}\rfloor}^{\prime},\cdots,C_{r-2}^{\prime}\rangle,$
which implies that
$\operatorname{Div}_{\operatorname{cusp}}^{0}(X_{0}(\mathfrak{n}/\mathfrak{p}))(K)/\beta_{\ast}(\operatorname{Div}_{\operatorname{cusp}}^{0}(X_{0}(\mathfrak{n}))(K))\cong(\mathbb{Z}/|\mathfrak{p}|\mathbb{Z})^{\lfloor\frac{r}{2}\rfloor-1}.$
6 Proof of Main Theorem
Fix a prime-power ideal $\mathfrak{n}=\mathfrak{p}^{r}\lhd A$ with $r\geq 2$. We prove Thm. 1.7 by induction on $r$. It is known that $\mathcal{C}(\mathfrak{p})=\mathcal{T}(\mathfrak{p})$. Assume that $\mathcal{C}(\mathfrak{p}^{r-1})_{\ell}=\mathcal{T}(\mathfrak{p}^{r-1})_{\ell}$ for a prime $\ell\nmid q(q-1)$. We want to show that $\mathcal{C}(\mathfrak{p}^{r})_{\ell}=\mathcal{T}(\mathfrak{p}^{r})_{\ell}$. Consider the following diagram:
[…]
We call $U_{\mathfrak{p}}(\mathcal{T}(\mathfrak{p}^{r}))_{\ell}\subset\mathcal{T}(\mathfrak{p}^{r})_{\ell}$ the old part of $\mathcal{T}(\mathfrak{p}^{r})_{\ell}$. By Lem. 5.3 and induction hypothesis, we have
$(\beta_{\mathfrak{p}})_{\ast}(\mathcal{C}(\mathfrak{p}^{r}))_{\ell}=\mathcal{C}(\mathfrak{p}^{r-1})_{\ell}=\mathcal{T}(\mathfrak{p}^{r-1})_{\ell}\supseteq(\beta_{\mathfrak{p}})_{\ast}(\mathcal{T}(\mathfrak{p}^{r}))_{\ell}.$
Thus, we have $U_{\mathfrak{p}}(\mathcal{C}(\mathfrak{p}^{r}))_{\ell}\supseteq U_{\mathfrak{p}}(\mathcal{T}(\mathfrak{p}^{r}))_{\ell}$. Moreover, since $\mathcal{C}(\mathfrak{p}^{r})\subseteq\mathcal{T}(\mathfrak{p}^{r})$, we have $U_{\mathfrak{p}}(\mathcal{C}(\mathfrak{p}^{r}))_{\ell}=U_{\mathfrak{p}}(\mathcal{T}(\mathfrak{p}^{r}))_{\ell}$. Note that for any prime $\ell\neq p$,
$U_{\mathfrak{p}}(\mathcal{C}(\mathfrak{p}^{r}))_{\ell}\cong\mathcal{C}(\mathfrak{p}^{r-1})_{\ell}\cong\left(\frac{\mathbb{Z}_{\ell}}{M(\mathfrak{p})\mathbb{Z}_{\ell}}\right)^{r-2}\times\frac{\mathbb{Z}_{\ell}}{N(\mathfrak{p})\mathbb{Z}_{\ell}},$
where $M(\mathfrak{p}):=\frac{|\mathfrak{p}|^{2}-1}{q^{2}-1}$ and
$N(\mathfrak{p}):=\begin{cases}\frac{|\mathfrak{p}|-1}{q^{2}-1},&\text{if $\deg(\mathfrak{p})$ is even.}\\ \frac{|\mathfrak{p}|-1}{q-1},&\text{otherwise.}\end{cases}$
Now, we look at the new part $\mathcal{T}(\mathfrak{p}^{r})_{\ell}[U_{\mathfrak{p}}]$ of $\mathcal{T}(\mathfrak{p}^{r})_{\ell}$. Recall that $(P(\mathfrak{n})_{\mathfrak{p}})$ is the sum of all the cusps of $X_{0}(\mathfrak{n})$ of height $\mathfrak{p}$. Define $\overline{C^{\prime}}\in\mathcal{C}(\mathfrak{p}^{r})$ by
$C^{\prime}:=\frac{|\mathfrak{p}|-1}{q-1}\cdot[0]-(P(\mathfrak{n})_{\mathfrak{p}})\in\operatorname{Div}_{\operatorname{cusp}}^{0}(X_{0}(\mathfrak{n}))(K).$ (4)
[…]
$g(C^{\prime})=\left(\left(\frac{\Delta}{\Delta_{\mathfrak{p}}}\right)^{|\mathfrak{p}|}\left(\frac{\Delta_{\mathfrak{p}^{2}}}{\Delta_{\mathfrak{p}}}\right)\right)\otimes\frac{1}{(q-1)(|\mathfrak{p}|^{2}-1)|\mathfrak{p}|^{r-2}}.$ (5)
From the above and [8, Lem. 3.6], the order of $\overline{C^{\prime}}$ in $\mathcal{C}(\mathfrak{p}^{r})$ divides $M(\mathfrak{p})|\mathfrak{p}|^{r-2}$. Since $\frac{\Delta_{\mathfrak{p}^{2}}}{\Delta_{\mathfrak{p}}}$ does not have a $p$-th root in $\mathcal{O}(\Omega)^{\ast}$ by [5, Cor. 3.18], we see that $|\mathfrak{p}|^{r-2}$ divides $\operatorname{ord}(\overline{C^{\prime}})$; cf. [8, Lem. 3.3]. Let $e:=\begin{pmatrix}\pi_{\infty}^{2}&\pi_{\infty}\\ 0&1\end{pmatrix}\in E(\mathscr{T})$; cf. [8, Sec. 2.2]. By [8, Lem. 2.4], we compute that
$|\mathfrak{p}|^{r-2}\cdot(\widetilde{r}\circ g(C^{\prime}))(e)=-\frac{|\mathfrak{p}|}{q}\cdot\frac{1}{M(\mathfrak{p})},$
which implies that $M(\mathfrak{p})$ also divides $\operatorname{ord}(\overline{C^{\prime}})$; cf. [8, Lem. 3.3]. Then we conclude that $\operatorname{ord}(\overline{C^{\prime}})=M(\mathfrak{p})|\mathfrak{p}|^{r-2}$.
Remark 6.1.
Combining the above with Thm. 1.4 and Lem. 5.3, the prime-to-$p$ part of $\mathcal{C}(\mathfrak{p}^{r})[U_{\mathfrak{p}}]$ is cyclic and generated by $|\mathfrak{p}|^{r-2}\cdot\overline{C^{\prime}}$ of order $M(\mathfrak{p})$. In contrast, the $p$-part of $\mathcal{C}(\mathfrak{p}^{r})[U_{\mathfrak{p}}]$ may not be cyclic.
Recall the following:
Lemma 6.2 ([12, Lem. 7.2]).
Suppose that $\ell$ is a prime not dividing $q(q-1)$. There is a natural injective homomorphism $\mathcal{T}(\mathfrak{n})_{\ell}\hookrightarrow\mathcal{E}_{00}(\mathfrak{n},\mathbb{Z}/\ell^{k}\mathbb{Z})$ for any $k\in\mathbb{Z}_{\geq 0}$ with $\ell^{k}\geq\#(\Phi_{\infty,\ell})$.
Proof..
The idea in the proof of [12, Lem. 7.2] is
$\mathcal{T}(\mathfrak{n})_{\ell}\hookrightarrow(\Phi_{\infty})_{\ell}[\mathfrak{E}(\mathfrak{n})]\hookrightarrow\mathcal{E}_{00}(\mathfrak{n},\mathbb{Z}/\ell^{k}\mathbb{Z})$
when $\ell^{k}\cdot(\Phi_{\infty})_{\ell}=0$ and $\ell\nmid q(q-1)$. ∎
From the above lemma, we obtain an injective $\mathbb{T}(\mathfrak{n})$-equivariant map:
$\frac{\mathbb{Z}_{\ell}}{M(\mathfrak{p})\mathbb{Z}_{\ell}}\cong\langle\overline{C^{\prime}}\rangle_{\ell}\subseteq\mathcal{C}(\mathfrak{n})_{\ell}[U_{\mathfrak{p}}]\subseteq\mathcal{T}(\mathfrak{n})_{\ell}[U_{\mathfrak{p}}]\hookrightarrow\mathcal{E}_{00}(\mathfrak{n},\mathbb{Z}/\ell^{k}\mathbb{Z})[U_{\mathfrak{p}}]$ (6)
for a prime $\ell\nmid q(q-1)$ and large enough $k$. It remains to show that
$\mathcal{E}_{00}(\mathfrak{p}^{r},\mathbb{Z}/\ell^{k}\mathbb{Z})[U_{\mathfrak{p}}]\hookrightarrow\frac{\mathbb{Z}_{\ell}}{M(\mathfrak{p})\mathbb{Z}_{\ell}}.$
Assume the above, then we have
$\mathcal{C}(\mathfrak{p}^{r})_{\ell}[U_{\mathfrak{p}}]=\mathcal{T}(\mathfrak{p}^{r})_{\ell}[U_{\mathfrak{p}}],$
which is the new part of $\mathcal{T}(\mathfrak{p}^{r})_{\ell}$. Combining the new and the old parts of $\mathcal{T}(\mathfrak{p}^{r})_{\ell}$, we obtain that $\mathcal{C}(\mathfrak{p}^{r})_{\ell}=\mathcal{T}(\mathfrak{p}^{r})_{\ell}$ for any prime $\ell\nmid q(q-1)$.
Define
$E_{\mathfrak{n}}:=\frac{1}{(q-1)(q^{2}-1)}\cdot\widetilde{r}\left(\left(\frac{\Delta}{\Delta_{\mathfrak{p}}}\right)^{|\mathfrak{p}|}\left(\frac{\Delta_{\mathfrak{p}^{2}}}{\Delta_{\mathfrak{p}}}\right)\right)\in\mathcal{H}(\mathfrak{n},\mathbb{Q}).$ (7)
Recall that the order of $\overline{C^{\prime}}$ in $\mathcal{C}(\mathfrak{p}^{r})$ is $M(\mathfrak{p})|\mathfrak{p}|^{r-2}$. Then by Thm. 3.4 and Eq. (5), we have $E_{\mathfrak{n}}=M(\mathfrak{p})|\mathfrak{p}|^{r-2}\cdot\widetilde{r}\circ g(C^{\prime})\in\mathcal{H}(\mathfrak{n},\mathbb{Z})$.
Lemma 6.3.
$E_{\mathfrak{n}}\in\mathcal{E}(\mathfrak{n},\mathbb{Z})[U_{\mathfrak{p}}]$.
Proof..
By Lem. 4.3 and 4.4, for any prime $\mathfrak{q}\lhd A$ not equal to $\mathfrak{p}$, we have $E_{\mathfrak{n}}|T_{\mathfrak{q}}=(|\mathfrak{q}|+1)E_{\mathfrak{n}}$ and $E_{\mathfrak{n}}|U_{\mathfrak{p}}=0$. Hence, $E_{\mathfrak{n}}$ is annihilated by both the actions of the Eisenstein ideal $\mathfrak{E}(\mathfrak{n})$ and the Hecke operator $U_{\mathfrak{p}}$. ∎
From now on, we always assume that $R$ is a coefficient ring. By composing with the map $\mathbb{Z}\rightarrow R$ sending $z$ to $z\cdot 1$, we also consider $E_{\mathfrak{n}}$ as an element of $\mathcal{E}(\mathfrak{n},R)[U_{\mathfrak{p}}]$. We claim that $E_{\mathfrak{n}}$ generates $\mathcal{E}(\mathfrak{n},R)[U_{\mathfrak{p}}]$ over $R$. Before we prove the claim, we recall the following:
Lemma 6.4 ([14, Lem. 2.15]).
Assume that $f\in\mathcal{H}(\mathfrak{n},R)$ is an eigenfunction of all $T_{\mathfrak{q}}$, $\mathfrak{q}\nmid\mathfrak{n}$; that is, $f|T_{\mathfrak{q}}=\lambda_{\mathfrak{q}}f$ for some $\lambda_{\mathfrak{q}}\in R$. Then the Fourier coefficients $f^{\ast}(\mathfrak{m})$, with $\mathfrak{m}$ coprime to $\mathfrak{n}$, are uniquely determined by $f^{\ast}(1)$ and the eigenvalues $\lambda_{\mathfrak{q}}$, $\mathfrak{q}\nmid\mathfrak{n}$.
Proof..
This follows from Prop. 4.2 and the fact that $(f|T_{\mathfrak{m}})^{\ast}(1)=|\mathfrak{m}|f^{\ast}(\mathfrak{m})$; cf. [11, Lem. 3.2]. ∎
We provide an Atkin-Lehner type result similar to [14, Thm. 2.18]:
Theorem 6.5.
Let $\mathfrak{n}=\mathfrak{p}^{r}$ with a prime $\mathfrak{p}\lhd A$ and $r\geq 2$. Suppose that $f\in\mathcal{H}(\mathfrak{n},R)$ is such that $f^{\ast}(\mathfrak{m})=0$ unless $\mathfrak{p}$ divides $\mathfrak{m}$. Then there exists $h\in\mathcal{H}(\mathfrak{n}/\mathfrak{p},R)$ such that $f=h|B_{\mathfrak{p}}$.
Proof..
Let $h:=f|B_{\mathfrak{p}}^{-1}$ and
$\Gamma_{0}(\mathfrak{n}/\mathfrak{p},\mathfrak{p}):=\left\{\begin{pmatrix}a&b\\ c&d\end{pmatrix}\in G(A)\mid c\in\mathfrak{n}/\mathfrak{p},~{}b\in\mathfrak{p}\right\}.$
By [14, Lem. 2.13], $h$ is $\Gamma_{0}(\mathfrak{n}/\mathfrak{p},\mathfrak{p})$-invariant and harmonic. Moreover, by Thm. 3.3, for any $e=\begin{pmatrix}\pi_{\infty}^{k}&y\\ 0&1\end{pmatrix}\in E(\mathscr{T})$ and $a\in A$, we have
$h\left(\begin{pmatrix}1&a\\ 0&1\end{pmatrix}e\right)=f\left(\begin{pmatrix}\pi_{\infty}^{k+\deg\mathfrak{p}}&y+\frac{a}{\mathfrak{p}}\\ 0&1\end{pmatrix}\right)$
$=f^{0}(\pi_{\infty}^{k+\deg\mathfrak{p}})+\sum_{\begin{subarray}{c}0\neq m\in A\\ \deg m\leq k+\deg\mathfrak{p}-2\end{subarray}}f^{\ast}(\operatorname{div}(m)\infty^{k+\deg\mathfrak{p}-2})\eta\left(my+\frac{ma}{\mathfrak{p}}\right)$
$=f^{0}(\pi_{\infty}^{k+\deg\mathfrak{p}})+\sum_{\begin{subarray}{c}0\neq m\in A\\ \deg m\leq k+\deg\mathfrak{p}-2\end{subarray}}f^{\ast}(\operatorname{div}(m)\infty^{k+\deg\mathfrak{p}-2})\eta(my)=h(e),$
where the third equality comes from the assumption that $f^{\ast}(\mathfrak{m})=0$ unless $\mathfrak{p}$ divides $\mathfrak{m}$. Thus, $h$ is also $\Gamma_{\infty}$-invariant, which completes the proof since $\Gamma_{0}(\mathfrak{n}/\mathfrak{p})$ is generated by $\Gamma_{0}(\mathfrak{n}/\mathfrak{p},\mathfrak{p})$ and $\Gamma_{\infty}$. ∎
Now, we are able to prove the following:
Proposition 6.6.
$\mathcal{E}(\mathfrak{n},R)[U_{\mathfrak{p}}]=R\cdot E_{\mathfrak{n}}$.
Proof..
Note that $R\cdot E_{\mathfrak{n}}\subseteq\mathcal{E}(\mathfrak{n},R)[U_{\mathfrak{p}}]$. It suffices to prove the reverse inclusion. By Eq. (2) and Lem. 3.6, we have
$E_{\mathfrak{n}}^{0}(\pi_{\infty}^{k})=0\text{~{}and~{}}E_{\mathfrak{n}}^{\ast}(1)=\frac{|\mathfrak{p}|}{q}\in R^{\ast}.$
Let $f\in\mathcal{E}(\mathfrak{n},R)[U_{\mathfrak{p}}]$, then consider
$\tilde{f}:=f-f^{\ast}(1)\cdot\frac{q}{|\mathfrak{p}|}\cdot E_{\mathfrak{n}}\in\mathcal{E}(\mathfrak{n},R)[U_{\mathfrak{p}}].$
First, by Lem. 6.4 and Thm. 6.5, since $\tilde{f}^{\ast}(1)=0$, there exists $h\in\mathcal{H}(\mathfrak{n}/\mathfrak{p},R)$ which is $\Gamma_{\infty}$-invariant such that $h|B_{\mathfrak{p}}=\tilde{f}$. Note that
$(h|B_{\mathfrak{p}})|U_{\mathfrak{p}}=\tilde{f}|U_{\mathfrak{p}}=0.$
Second, by [12, Lem. 2.23], since $h$ is $\Gamma_{\infty}$-invariant, we have
$|\mathfrak{p}|\cdot h=(h|B_{\mathfrak{p}})|U_{\mathfrak{p}}=0.$
Hence, $h=0$ and $\tilde{f}=h|B_{\mathfrak{p}}=0$, which implies that
$f=f^{\ast}(1)\cdot\frac{q}{|\mathfrak{p}|}E_{\mathfrak{n}}\in R\cdot E_{\mathfrak{n}}.$
Note that
$\begin{pmatrix}1&0\\ \mathfrak{p}&1\end{pmatrix}\cdot[\infty]:=\begin{pmatrix}1&0\\ \mathfrak{p}&1\end{pmatrix}\cdot\begin{bmatrix}1\\ \mathfrak{n}\end{bmatrix}=\begin{bmatrix}1\\ \mathfrak{p}\end{bmatrix}$
on the cusps of $X_{0}(\mathfrak{n})$. Let $e_{j}:=\begin{pmatrix}1&0\\ \mathfrak{p}&1\end{pmatrix}\begin{pmatrix}\pi_{\infty}^{-j}&0\\ 0&1\end{pmatrix}\in E(\mathscr{T})$ with $j\geq 0$. Then $\{\Gamma_{0}(\mathfrak{n})\cdot e_{j}\mid j\geq 0\}$ forms an end on the quotient graph of the Bruhat-Tits tree $\Gamma_{0}(\mathfrak{n})\backslash\mathscr{T}$ corresponding to the cusp $\begin{bmatrix}1\\ \mathfrak{p}\end{bmatrix}$ on $X_{0}(\mathfrak{n})$; cf. [14, Def. 2.5] and [5, p. 184]. Before computing $\mathcal{E}_{0}(\mathfrak{n},R)[U_{\mathfrak{p}}]$, we consider the following lemma:
Lemma 6.7 (cf. [8, Lem. 2.4]).
Let $j\geq 0$. Then we have
  1. 1.
    $\widetilde{r}(\Delta)(\begin{pmatrix}1&0\\ \mathfrak{p}&1\end{pmatrix}\begin{pmatrix}\pi_{\infty}^{-j}&0\\ 0&1\end{pmatrix})=-(q-1)(q^{j+1}-q-1).$
  2. 2.
    Assume that $k\geq 1$ and $j\geq\max\{(k-2)\deg(\mathfrak{p}),0\}$. Then
    $\widetilde{r}(\Delta_{\mathfrak{p}^{k}})(\begin{pmatrix}1&0\\ \mathfrak{p}&1\end{pmatrix}\begin{pmatrix}\pi_{\infty}^{-j}&0\\ 0&1\end{pmatrix})=-(q-1)(q^{j+1}|\mathfrak{p}|^{2-k}-q-1).$
Proof..
For $k\geq 0$ and $j\geq 0$, we have
$\begin{pmatrix}\mathfrak{p}^{k}\pi_{\infty}^{-j}&0\\ \mathfrak{p}\pi_{\infty}^{-j}&1\end{pmatrix}\begin{pmatrix}1&\mathfrak{p}^{1-k}\pi_{\infty}^{j+(2-k)\deg(\mathfrak{p})}\\ 0&-(\mathfrak{p}\pi_{\infty}^{\deg(\mathfrak{p})})^{2-k}\end{pmatrix}\begin{pmatrix}\mathfrak{p}^{-1}\pi_{\infty}^{j+1}&0\\ 0&\mathfrak{p}^{-1}\pi_{\infty}^{j+1}\end{pmatrix}$
$=\begin{pmatrix}\mathfrak{p}^{k-1}\pi_{\infty}&\pi_{\infty}^{j+1+(2-k)\deg(\mathfrak{p})}\\ \pi_{\infty}&0\end{pmatrix}=\begin{pmatrix}\pi_{\infty}^{j+1+(2-k)\deg(\mathfrak{p})}&\mathfrak{p}^{k-1}\\ 0&1\end{pmatrix}\begin{pmatrix}0&1\\ \pi_{\infty}&0\end{pmatrix}.$
Thus, $\begin{pmatrix}\mathfrak{p}^{k}\pi_{\infty}^{-j}&0\\ \mathfrak{p}\pi_{\infty}^{-j}&1\end{pmatrix}=\begin{pmatrix}\pi_{\infty}^{j+1+(2-k)\deg(\mathfrak{p})}&\mathfrak{p}^{k-1}\\ 0&1\end{pmatrix}\begin{pmatrix}0&1\\ \pi_{\infty}&0\end{pmatrix}$ in $E(\mathscr{T})$.
First, assume that $k=0$ and $j\geq 0$, then
$\begin{pmatrix}\pi_{\infty}^{j+1+2\deg(\mathfrak{p})}&\mathfrak{p}^{-1}\\ 0&1\end{pmatrix}\begin{pmatrix}0&1\\ \pi_{\infty}&0\end{pmatrix}\begin{pmatrix}\mathfrak{p}\pi_{\infty}^{\deg(\mathfrak{p})}&\pi_{\infty}^{j+\deg(\mathfrak{p})}\\ 0&-(\mathfrak{p}\pi_{\infty}^{\deg(\mathfrak{p})})^{-1}\end{pmatrix}\begin{pmatrix}\pi_{\infty}^{-\deg(\mathfrak{p})}&0\\ 0&\pi_{\infty}^{-\deg(\mathfrak{p})}\end{pmatrix}$
$=\begin{pmatrix}\pi_{\infty}&0\\ \mathfrak{p}\pi_{\infty}&\pi_{\infty}^{j+1}\end{pmatrix}=\begin{pmatrix}0&1\\ 1&\mathfrak{p}\end{pmatrix}\begin{pmatrix}\pi_{\infty}^{j+1}&0\\ 0&1\end{pmatrix}\begin{pmatrix}0&1\\ \pi_{\infty}&0\end{pmatrix}.$
This implies that
$\widetilde{r}(\Delta)(\begin{pmatrix}1&0\\ \mathfrak{p}&1\end{pmatrix}\begin{pmatrix}\pi_{\infty}^{-j}&0\\ 0&1\end{pmatrix})=\widetilde{r}(\Delta)(\begin{pmatrix}\pi_{\infty}^{-j}&0\\ \mathfrak{p}\pi_{\infty}^{-j}&1\end{pmatrix})$
$=\widetilde{r}(\Delta)(\begin{pmatrix}\pi_{\infty}^{j+1+2\deg(\mathfrak{p})}&\mathfrak{p}^{-1}\\ 0&1\end{pmatrix}\begin{pmatrix}0&1\\ \pi_{\infty}&0\end{pmatrix})$
$=\widetilde{r}(\Delta)(\begin{pmatrix}0&1\\ 1&\mathfrak{p}\end{pmatrix}\begin{pmatrix}\pi_{\infty}^{j+1}&0\\ 0&1\end{pmatrix}\begin{pmatrix}0&1\\ \pi_{\infty}&0\end{pmatrix})$
$=-\widetilde{r}(\Delta)(\begin{pmatrix}\pi_{\infty}^{j+1}&0\\ 0&1\end{pmatrix})=-(q-1)(q^{j+1}-q-1),$
where the penultimate equality comes from [5, Eq. (2.3)].
Second, assume that $k\geq 1$ and $j\geq\max\{(k-2)\deg(\mathfrak{p}),0\}$, then
$\widetilde{r}(\Delta_{\mathfrak{p}^{k}})(\begin{pmatrix}1&0\\ \mathfrak{p}&1\end{pmatrix}\begin{pmatrix}\pi_{\infty}^{-j}&0\\ 0&1\end{pmatrix})$
$=\widetilde{r}(\Delta)(\begin{pmatrix}\mathfrak{p}^{k}&0\\ 0&1\end{pmatrix}\begin{pmatrix}1&0\\ \mathfrak{p}&1\end{pmatrix}\begin{pmatrix}\pi_{\infty}^{-j}&0\\ 0&1\end{pmatrix})=\widetilde{r}(\Delta)(\begin{pmatrix}\mathfrak{p}^{k}\pi_{\infty}^{-j}&0\\ \mathfrak{p}\pi_{\infty}^{-j}&1\end{pmatrix})$
$=\widetilde{r}(\Delta)(\begin{pmatrix}\pi_{\infty}^{j+1+(2-k)\deg(\mathfrak{p})}&\mathfrak{p}^{k-1}\\ 0&1\end{pmatrix}\begin{pmatrix}0&1\\ \pi_{\infty}&0\end{pmatrix})$
$=\widetilde{r}(\Delta)(\begin{pmatrix}1&\mathfrak{p}^{k-1}\\ 0&1\end{pmatrix}\begin{pmatrix}\pi_{\infty}^{j+1+(2-k)\deg(\mathfrak{p})}&0\\ 0&1\end{pmatrix}\begin{pmatrix}0&1\\ \pi_{\infty}&0\end{pmatrix})$
$=-\widetilde{r}(\Delta)(\begin{pmatrix}\pi_{\infty}^{j+1+(2-k)\deg(\mathfrak{p})}&0\\ 0&1\end{pmatrix})=-(q-1)(q^{j+1}|\mathfrak{p}|^{2-k}-q-1),$
where the penultimate equality again comes from [5, Eq. (2.3)]. ∎
Lemma 6.8.
Let $R\left[M(\mathfrak{p})\right]:=\{c\in R\mid M(\mathfrak{p})\cdot c=0\}$. Then
$\mathcal{E}_{0}(\mathfrak{n},R)[U_{\mathfrak{p}}]\subseteq R\left[M(\mathfrak{p})\right]\cdot E_{\mathfrak{n}}.$
Proof..
By Prop. 6.6, $\mathcal{E}_{0}(\mathfrak{n},R)[U_{\mathfrak{p}}]$ is contained in $R\cdot E_{\mathfrak{n}}$. Suppose that $c\cdot E_{\mathfrak{n}}\in\mathcal{E}_{0}(\mathfrak{n},R)[U_{\mathfrak{p}}]$ for some $c\in R$. Let $e_{j}:=\begin{pmatrix}1&0\\ \mathfrak{p}&1\end{pmatrix}\begin{pmatrix}\pi_{\infty}^{-j}&0\\ 0&1\end{pmatrix}\in E(\mathscr{T})$ with $j\geq 0$. By Lem. 6.7, we have $c\cdot E_{\mathfrak{n}}(e_{j})=c\cdot q^{j+1}M(\mathfrak{p})$ for $j\geq 0$. Since $q\in R^{\ast}$, we have $c\cdot E_{\mathfrak{n}}(e_{j})=0$ for any $j\gg 0$ if and only if $c\in R\left[M(\mathfrak{p})\right]$, which completes the proof. ∎
Now, we prove the following:
Proposition 6.9.
Let $\mathfrak{n}=\mathfrak{p}^{r}\lhd A$ with a prime $\mathfrak{p}\lhd A$ and $r\geq 2$. Assume that $\ell$ is a prime not dividing $q(q-1)$. Then
$C(\mathfrak{n})_{\ell}[U_{\mathfrak{p}}]=\mathcal{T}(\mathfrak{n})_{\ell}[U_{\mathfrak{p}}].$
Proof..
Let $R:=\mathbb{Z}/\ell^{k}\mathbb{Z}$ with $k\gg 0$ such that $\mathcal{T}(\mathfrak{n})_{\ell}\hookrightarrow\mathcal{E}_{00}(\mathfrak{n},R)$ in Lem. 6.2. By Lem. 6.8,
$\mathcal{E}_{00}(\mathfrak{n},R)[U_{\mathfrak{p}}]\subseteq\mathcal{E}_{0}(\mathfrak{n},R)[U_{\mathfrak{p}}]\subseteq R\left[M(\mathfrak{p})\right]\cdot E_{\mathfrak{n}}\cong\frac{\mathbb{Z}_{\ell}}{M(\mathfrak{p})\mathbb{Z}_{\ell}}.$
Thus, the equality holds in Sequence (6). In particular, we have
$\mathcal{C}(\mathfrak{n})_{\ell}[U_{\mathfrak{p}}]=\mathcal{T}(\mathfrak{n})_{\ell}[U_{\mathfrak{p}}].$
From the above, we obtain Main Theorem.